# Quantifying Privacy Leakage in Split Inference via Fisher-Approximated Shannon Information Analysis

### v1

@@ 4.3 Computing Privacy Leakage
Fisher information is a fundamental concept in statistics hypothesis testing (Martens, 2020), which is related to the asymptotic variability of an estimator, where higher FI means lower estimation error:
Definition 4.5.
(Fisher Information Matrix) (Cover, 1999). For two random variables $x\sim p(x)$ and $z\sim p(z|x)$, the element of Fisher Information Matrix (FIM) or the Fisher Information (FI) $F_{z|x}$ of $z$ w.r.t. the $x$ is defined as:
$F_{z|x}(i,j)$ $=\mathbb{E}_{z}[(\frac{\partial}{\partial x_{i}}\log p(z|x))(\frac{\partial}{\partial x_{j}}\log p(z|x))]$ (12)
$=-\mathbb{E}_{z}[\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\log p(z|x)],$ (13)
where the two formulations are equivalent (Martens, 2020) and $\log p(z|x)$ is the log conditional probability density function (p.d.f.) of $z$ given $x$.
Note that, FI is defined based on the distribution; however, in the SI, the neural network $f_{\theta 1}$ is a deterministic function, and $p(z|x)=p(f_{\theta 1}(x)|x)=1$. Therefore, we fit FI into SI system by following (Maeng et al., 2023; Martens, 2020) which combine $f_{\theta 1}$ with a Gaussian distribution: $\tilde{f}_{\theta 1}(\cdot)=r(f_{\theta 1}(\cdot))$, where $r(z)=z+\delta,\delta\sim\mathcal{N}(0,\sigma^{2})$. To this end, we can calculate FI of raw input $x$ contained by the smashed data $z$ as:
@@
Theorem 4.6.
(Cramér-Rao Bound) (Cover, 1999). The mean-squared error of any unbiased estimator $T(z)$ of the parameter $x$ is lower bounded by the reciprocal of the Fisher information:
@@
Cramér-Rao Bound gives the covariance’s lower bound of the estimator $T(z)$, or $\hat{X}$ given $X=x$ in our DRA settings of SI, which can only be achieved by the “strongest” DRA adversary $\mathcal{A}^{s}$ (see Section 4.2) who has the lowest estimation error. In other words, given $X=x$, the reconstructed distribution $p(\hat{X}^{s}|X=x)$ of the unbiased $\mathcal{A}^{s}$ has a covariance of $Cov(\hat{X}^{s})=F_{z|x}^{-1}$ and the mean of $x$.
Entropy Maximization. Among all the distributions with the same mean and covariance, the normal distribution $\mathcal{N}(x,F_{z|x}^{-1})$ maximizes entropy (Tan et al., 2024). To obtain the supremum of $H(\hat{X}^{s}|X=x)$ we let $p(\hat{X}^{s}|X=x)=\mathcal{N}(x,F_{z|x}^{-1})$ and further compute the $H(\hat{X}^{s}|X=x)$ as
$H(\hat{X}^{s}|X=x)=\frac{1}{2}\log(\frac{(2\pi e)^{d_{x}}}{det(F_{z|x})}),$ (16)
where the $d_{x}$ is the dimension of raw input $x$ and $det(F_{z|x})$ is the determinant of $F_{z|x}$.
Now, with the raw input X, we can directly calculate the approximation of privacy leakage against the strongest adversary $\mathcal{A}^{s}$ in Definition 4.3 as:
$\mathcal{L}^{\mathcal{A}^{s}}_{f_{\theta 1},X}=-H(\hat{X}^{s}|X)=-\mathbb{E}_{x\sim p(X)}[\frac{1}{2}\log(\frac{(2\pi e)^{d_{x}}}{det(F_{z|x})})],$ (17)
where all the terms can be easily implemented.
Acceleration. The calculation of $\mathcal{L}^{\mathcal{A}^{s}}$ involves computing the Jacobian of $z$ w.r.t. $x$, which is time-consuming when the $d_{x}$ gets large (e.g., for a $32\times 32\times 3$ image in CIFAR10 (Krizhevsky et al., 2009), the number of rows of $F_{z|x}$ is 3072). We speed up this process by first approximating $F_{z|x}$ with its diagonal version, which is widely used in transfer learning (Fan et al., 2021; Xue et al., 2021):
$\mathcal{L}^{\mathcal{A}^{s}}_{f_{\theta 1},X}=$ $-\frac{1}{2}\mathbb{E}_{x\sim p(X)}[d_{x}\log(2\pi e)-\log\det(F_{z|x})].$ (18)
Second, we use subsampling techniques such as average pooling to reduce input dimensions for image datasets.
Privacy Leakage per Dimension. Although $\mathcal{L}^{\mathcal{A}^{s}}$ can now outline the privacy leakage risk of an SI system, its value exhibits considerable variation w.r.t. the dimensionality of the input data, e.g., it tends to increase with the dimension of the input data. Therefore, we use the average FSInfo across all dimensions as a metric:
$\overline{\mathcal{L}^{\mathcal{A}^{s}}_{f_{\theta 1},X}}=\frac{\mathcal{L}^{\mathcal{A}^{s}}_{f_{\theta 1},X}}{d_{x}}.$ (19)
Definition 4.7.
@@
$FSInfo=$ $\overline{\mathcal{L}^{\mathcal{A}^{s}}_{f_{\theta 1},X}}$ (20)
$=$ $-\frac{1}{2d_{x}}\mathbb{E}_{x\sim p(X)}[d_{x}\log(2\pi e)-\log\det(F_{z|x})].$ (21)
@@
Definition 4.7 gives the essence link in DRA between the Fisher information and Shannon information, where the maximal Shannon information gain can be achieved by adversary $\mathcal{A}^{s}$ whose estimation error is lower bounded by the Fisher information.
FSInfo serves as a crucial metric for both regulatory compliance and system optimization. From a regulatory perspective, it enables effective supervision of ML products, preventing market entry for systems with excessive privacy leakage risks. Adversaries can leverage FSInfo to identify vulnerable SI systems, optimizing their attack strategies, while practitioners can utilize it to select architectures and methodologies that minimize privacy leakage, thereby enhancing overall system security.
5 Experiments
In this section, we first apply our FSInfo on several models and datasets and compare it with other privacy metrics to validate the efficacy of FSInfo as a metric and demonstrate that privacy leakage is common in different settings. Second, we experimentally analyze factors that influence privacy leakage. We also provided results for privacy leakage for defenses SI systems in Appendix D.
5.1 Setups
@@
5.1.1 Datasets and Models.
We follow the common settings of existing works on SI privacy attacks and defense (Yin et al., 2023; Noorbakhsh et al., 2024), where detailed dataset and model settings are provided in Appendix B. Image-based tasks are ubiquitous in ML applications, yet they present significant privacy concerns; hence, as a pioneering work, our first experimental study is primarily concentrated on image tasks. We use two datasets for our experiments: CIFAR-10 (Krizhevsky et al., 2009), and MNIST (LeCun et al., 1998). We use the train set for model training, and the test set for privacy metrics computing. We use four models including VGG-5 (Simonyan & Zisserman, 2014), VGG-9 (Simonyan & Zisserman, 2014), ResNet-18 (He et al., 2016), ResNet-34 (He et al., 2016). The number of critical layers and architecture of these models implemented for CIFAR-10 are presented in Appendix B. The split layer is identified by the index of that critical layer, starting from 1.
5.1.2 Compared Methods and Metrics
We empirically examined the FSInfo by comparing it with the three state-of-the-art (SOTA) methods: Diagonal Fisher information leakage (we use its reciprocal version: 1/dFIL) (Guo et al., 2022; Maeng et al., 2023), DLoss (Deng et al., 2024; Székely et al., 2007; Sun et al., 2021, 2022), and Mutual Information (MI) (Mireshghallah et al., 2020).
Unlike MIA settings where training data ratios in test sets provide measurable ground truth (Song & Mittal, 2021), there is no trivial way to experimentally control the ground truth privacy leakage due to the poor interpretability of neural networks (Deng et al., 2024). We therefore adopt reconstruction accuracy as our primary metric, aligning theoretical formulations with empirical observations. Following Section 3.2’s threat model, we adopt an NN-based DRA approach (He et al., 2019), as empirical evidence demonstrates superior reconstruction quality of NN-based methods compared to MLE-based approaches (Yin et al., 2023) (See Appendix A for detailed explanations). Evaluation incorporates standard DRA metrics: visual invertibility, Structural Similarity Index Measure (SSIM), and MSE. Note that these metrics only establish a lower bound for actual privacy leakage—at least what the attacker has already obtained has been leaked. The detailed introduction of the compared methods and metrics is given in Appendix C.
For comparison with SOTA privacy quantification methods with different value spans, inspired by Tan et al. (Tan et al., 2024), we construct the improvement criteria. Specifically, with privacy metrics denoted by $Metric$, the improvement (IMP) of compared methods and MSE is:
$IMP\ (Metric)$ $=({Metric_{b}}-{Metric})/{Metric_{b}}$ (22)
$IMP\ (MSE)$ $=(\frac{1}{MSE_{b}}-\frac{1}{MSE})/\frac{1}{MSE_{b}},$ (23)
where the experimental result with the highest degree of privacy leakage is designated with subscript $b$.
5.2 Validation of FSInfo
5.2.1 Validation by Attack Success Metrics
Figure 4: SSIM vs. FSInfo on three different models and CIFAR-10.
Figure 5: MSE vs. FSInfo on three different models and CIFAR-10.
Figure 6: Evaluating different metrics in the Metric-MSE plane for four different split points of VGG-5 on CIFAR-10. We also report the Pearson correlation coefficient of each metric.
Figure 7: Evaluating different metrics in the Metric-MSE plane for four different split points of ResNet-18 on CIFAR-10. We also report the Pearson correlation coefficient of each metric.
We first validate the efficacy of the FSInfo by comparing it with experimental DRA attack results. Figure 3 visualizes the raw input and the reconstructed image along with the averaged layer FSInfo. Generally, we observe that as the image clarity decreases, the corresponding FSInfo also decreases. This indicates our FSInfo has a strong correlation with the visual quality of reconstructed images.
Then, we compare FSInfo in Figure 4 and Figure 5 with SSIM and MSE, two widely adopted objective metrics for DRA success. First, We can see that FSInfo exhibits a positive correlation with SSIM and a negative correlation with MSE as expected, validating that the FSInfo reveals privacy leakage correctly. Additionally, there is a rise in the line’s gradient of ResNet-18 when FSInfo approaches 0.5 in Figure 4. This demonstrates that chances are the adversary captures less privacy leakage than the actual extent, or the adversary’s capabilities are insufficient to exploit the full extent of the leaked privacy. This is also the reason for the downward trend of VGG-9’s line in Figure 4.
5.2.2 Comparing with Other Methods
Theoretically, FSInfo provides key theoretical advantages over alternatives: MI overlooks DRA-specific characteristics, while 1/dFIL’s MSE-only lower bound lacks generalizability. As demonstrated in Table 1 of the manuscript, FSInfo addresses both limitations.
However, experimentally, there is no ground truth for comparing the accuracy of different privacy metrics using diverse theoretical language to define privacy leakage, and thus, previous works like 1/dFIL (Guo et al., 2022) don’t compare their methods with others on the “accuracy” of quantified leakage. Instead, they only show that their metric is related to the attack success rate in the trend. Therefore, we compare our FSInfo with 1/dFIL, DLoss, and MI in the Metric-MSE plane and calculate their Pearson correlation coefficients (PCC) to showcase that the FSInfo has a stronger correlation with the changes of the adversary’s MSE. As is shown in Figure 6 and Figure 7, the FSInfo achieves the best alignment with empirical observations among all metrics, whose PPC are both 0.99 on VGG-5 and ResNet-18. It also signifies that different privacy metrics themselves can disagree with each other on their way to approaching the ground-truth privacy leakage.
5.3 Factors Affecting Privacy Leakage
5.3.1 Intrinsic Data Characteristics

### v2

@@ IV-C Computing Privacy Leakage
Fisher information is a fundamental concept in statistics hypothesis testing [martens2020new], which is related to the asymptotic variability of an estimator, where higher FI means lower estimation error:
Definition IV.5.
(Fisher Information Matrix) [cover1999elements]. For two random variables $x\sim p(x)$ and $z\sim p(z|x)$, the element of Fisher Information Matrix (FIM) or the Fisher Information (FI) $F_{z|x}$ of $z$ w.r.t. the $x$ is defined as:
$F_{z|x}(i,j)$ $=\mathbb{E}_{z}[(\frac{\partial}{\partial x_{i}}\log p(z|x))(\frac{\partial}{\partial x_{j}}\log p(z|x))],$ (9)
where $\log p(z|x)$ is the log conditional probability density function (p.d.f.) of $z$ given $x$.
Note that, FI is defined based on the distribution; however, in the SI, the neural network $f_{\theta 1}$ is a deterministic function, and $p(z|x)=p(f_{\theta 1}(x)|x)=1$. Therefore, we follow [maeng2023bounding, martens2020new] to fit FI into DL models by combining $f_{\theta 1}$ with a random noise: $\tilde{f}_{\theta 1}(\cdot)=r(f_{\theta 1}(\cdot))$, where $r(z)=z+\delta,\delta\sim\mathcal{N}(0,\sigma^{2})$. To this end, we can calculate FI of raw input $x$ contained by the smashed data $z$ as:
@@
Theorem IV.6.
(Cramér-Rao Bound) [cover1999elements]. The mean-squared error of any unbiased estimator $T(z)$ of the parameter $x$ is lower bounded by the reciprocal of the Fisher information:
@@
Theorem IV.6 establishes the optimality of the SA. Considering the DRA adversary as an estimator $T(z)$, the strongest adversary should have the minimal covariance $F_{Z|X}^{-1}$ and a mean of $\mathbb{E}[X]$ (according to the unbiased assumptions in Section III-B). The $H(\hat{X}|X)$, no matter what types of distribution the $p(\hat{X}|X)$ and p(X^)p($\hat${X}) are, is smaller than the entropy of a Gaussian distribution with the same covariance and mean. Because, with the same mean and covariance, the Gaussian distribution has the largest entropy [sec24-defending-FL-ita]. Therefore, the possibly least captured leakage of the SA is: $H(\mathcal{N}(x,F_{Z|X}^{-1}))=-\frac{1}{2}\log(\frac{(2\pi e)^{d_{x}}}{det(F_{Z|X})})$. Consequently, we summarize the above discussion into the following Theorem IV.7.
Theorem IV.7.
(Privacy Leakage Lower Bound of SA). A Lower Bound of the privacy leakage against the strongest adversary is:
$\mathcal{L}^{\mathcal{A}^{s}}_{f_{\theta 1},X}=-H(\hat{X}^{s}|X)\geq-\frac{1}{2}\log(\frac{(2\pi e)^{d_{x}}}{det(F_{Z|X})}),$ (12)
where the $det(F_{Z|X})$ is the determinant of $F_{Z|X}$.
Now, with the raw input $X$, we can directly calculate the lower bound of privacy leakage against the strongest adversary $\mathcal{A}^{s}$ in Definition IV.3 through Equation 12, where all the terms can be easily implemented.
Acceleration. The calculation of $\frac{1}{2}\log(\frac{(2\pi e)^{d_{x}}}{det(F_{Z|X})})$ involves computing the Jacobian of $z$ w.r.t. $x$, which is time-consuming when the $d_{x}$ gets large (e.g., for a $32\times 32\times 3$ image in CIFAR10 [krizhevsky2009learning], the number of rows of $F_{Z|X}$ is 3072). We speed up this process by first approximating $F_{Z|X}$ with its diagonal version, which is widely used in transfer learning [fan2021discriminative, xue2021toward]:
$-\frac{1}{2}\log(\frac{(2\pi e)^{d_{x}}}{det(F_{Z|X})})=$ $-\frac{1}{2}(d_{x}\log(2\pi e)-\sum_{i}^{d_{x}}\log(\Lambda_{i})),$ (13)
where the $\Lambda_{i}$ is the diagonal element of $F_{Z|X}$. Second, we use subsampling techniques such as random sampling and average pooling to reduce input dimensions $d_{x}$.
Privacy Leakage per Dimension. Although $\mathcal{L}^{\mathcal{A}^{s}}$ can now outline the privacy leakage risk of an SI system, its value exhibits considerable variation w.r.t. the dimensionality of the input data, e.g., it tends to increase with the dimension of the input data. Therefore, we adopt the average leakage across all dimensions. FSInfo is a computable privacy metric, defined as an unbiased estimate of this average leakage using data pairs $\{x,y\}\in\mathcal{U}$:
Definition IV.8.
@@
$FSInfo=$ $-\frac{1}{2d_{x}}\mathbb{E}_{x\sim p(X)}[d_{x}\log(2\pi e)-\sum_{i}^{d_{x}}\log(\lambda_{i})],$ (14)
where the $\lambda_{i}$ is the diagonal element of $F_{z|x}$.
@@
Definition IV.8 gives the essence link in DRA between the Fisher information and Shannon information, where the maximal Shannon information gain can be achieved by adversary $\mathcal{A}^{s}$ whose estimation error is lower bounded by the Fisher information.
IV-D Guiding Calculation of Defensive Noise
FSInfo not only quantifies privacy leakage in SI systems but also provides a risk signal that can guide the design of defense mechanisms. Based on this idea, we develop FSInfoGuard, a defense against DRAs for SI systems.
Privacy protection in split inference differs fundamentally from federated learning. In SI, smashed data preserves a one-to-one mapping to the raw input rather than aggregating information, meaning that batch size—an important factor in FL privacy—has limited influence on SI privacy leakage [geiping2020inverting]. Consequently, obfuscation-based defense mechanisms, particularly those utilizing regularization or closed-form noise injection techniques, have become predominant in split inference systems [singh2024simba]. Regularization methods optimize the client-side model with heuristic regularization terms in the loss function to guide it to output smashed data that is hard to invert. Closed-form defenses directly add noise with an analytical noise scale on the smashed data to cater for the desired security level (e.g., metric-DP [singh2023posthoc]).
These obfuscation-based defenses typically rely on a privacy-related objective or feedback signal to perturb the smashed data. Nopeek [ICDM-20-nopeek] and Shredder [mireshghallah2020shredder] are representative regularization-based approaches. Nopeek adopts adversarial representation learning to encourage the bottom model to produce smashed data containing less information about the raw input, using the loss: $\alpha DLoss(x,z)+CCE$, where $DLoss$ is a privacy metric estimates the correlation between $x$ and $z$ and $CCE$ is the cross-entropy loss. Shredder instead trains a noise layer to generate Gaussian perturbations that explicitly reduce the estimated mutual information (a privacy metric) between $x$ and $z$. inv_dFIL_def is a closed-form defense that adds Gaussian noise with scale $\sigma=\sqrt{\frac{\text{Tr}(J^{\top}_{f_{\theta 1}}J_{f_{\theta 1}})}{d\times\text{dFIL}}}$, where $f_{\theta 1}$ is the bottom model, $d$ is the input dimension, and dFIL is the target privacy metric. This noise guarantees the desired dFIL privacy level.
FSInfo can serve as a standalone privacy metric for designing both regularization-based and closed-form defenses. In principle, one could minimize FSInfo directly by adding it to the training loss; however, this empirical regularization strategy offers no provable security guarantees. Motivated by this limitation, we introduce FSInfoGuard, a closed-form noise-based defense that injects Gaussian noise calibrated to meet a target FSInfo level. From Definition IV.8 and Equation 10, achieving a specific FSInfo requires adding Gaussian noise with zero mean and scale $\sigma=\frac{\det(J_{f_{\theta 1}}^{T}J_{f_{\theta 1}})^{\frac{1}{2d}}}{e^{FSInfo}(2\pi e)^{\frac{1}{2}}}$ to the smashed data. FSInfoGuard applies this noise during both training and inference: during training, the model learns from perturbed smashed data, and during inference, the injected noise suppresses privacy leakage and mitigates DRAs.
V Experiments
In this section, we first introduce the experimental setups. Then, we evaluate FSInfo across various models, datasets, and defenses to validate its effectiveness and highlight the prevalence of privacy leakage. We then show how FSInfo can guide the design of privacy-preserving architectures and methods by comparing the privacy-utility tradeoff of FSInfoGuard with other defenses. Finally, we analyze three factors that influence privacy leakage: data characteristics, model size, and overfitting.
V-A Setups
V-A1 Datasets and Models
We use four image datasets (CelebA [liu2015faceattributes], CIFAR-10 [krizhevsky2009learning], TinyImageNet [le2015tiny], and MNIST [lecun1998gradient]), two tabular datasets (Purchase100 [sec21-evaluation-ml] and Home Credit [home-credit-kaggle]), and one text dataset (GLUE/SST2 [wang2018glue]) for our experiments. We use the train set for model training, and the test set for FSInfo computing. We use VGG-5 [simonyan2014very], VGG-9 [simonyan2014very], ResNet-18 [he2016deep], ResNet-34 [he2016deep] for image data, FCNs for tabular data, and DistilBert [sanh2019distilbert] for text data. The number of critical layers and the architecture of these models implemented are presented in the Appendix. The split layer is identified by the index of that critical layer, starting from 1. We only present the results for the image dataset for lack of space, and the additional results can be found in the Appendix.
V-A2 Attacks and Defenses
Following [noorbakhsh2024inf2guard], we adopt an NN-based DRA approach [blackbox]. We adopt three commonly used metrics to evaluate the attack performance: visual invertibility, Structural Similarity Index Measure (SSIM), and Mean Squared Error (MSE). Visual invertibility measures the similarity of the raw input and the reconstructed one by human perception [sun2024privacy]. We visualize the reconstructed images and observe their clarity and similarity to the raw input. This metric offers semantic judgment for privacy information observed by the adversary. A higher visual quality means a larger privacy leakage. SSIM is used to assess the perceived quality of generated images, which aligns closely with human visual perception. A higher SSIM indicates greater privacy leakage. MSE is also an objective criterion, which measures the error or distance between two images. A large MSE indicates less privacy leakage. Note that the attack efficacy only establishes a lower bound on the actual privacy leakage, rather than the ground truth—at least, what the attacker has already obtained has been leaked.
We use three defense methods in the experiments, including two regularization defenses (Nopeek [ICDM-20-nopeek] and Shredder [mireshghallah2020shredder]) and one closed-form defense (inv_dFIL_def [maeng2023bounding]). First, we adjust the defense strength of Nopeek and inv_dFIL_def to observe the corresponding FSInfo. Second, we compare the privacy-utility tradeoff between FSInfoGuard and these three defenses.
V-B Evaluating FSInfo among Different Datasets
(a) VGG-5, VGG-9, ResNet-18 on CIFAR-10.
(b) VGG-5 and VGG-9 on MNIST.
(c) DistilBert on GLUE/SST2.
(d) FCNs on Purchase100 and Home Credit.
Figure 3: MSE vs. FSInfo across different datasets and models.
In this section, we compute FSInfo on multiple modalities to examine whether it effectively reflects the degree of privacy leakage and aligns with the DRA adversary’s MSE. For each dataset, we vary the model’s split point to obtain smashed data with different exposure levels, enabling a clearer comparison between FSInfo and actual attack performance. For image (CIFAR-10, MNIST) and tabular (Purchase, Credit) datasets, we apply the inverse-network DRA and compute FSInfo of the smashed data with respect to the raw inputs. For text data, the discrete inputs are tokenized and passed through an embedding layer before entering the transformer. In this case, following Maeng et al. [maeng2023bounding], we adopt an MLE-based DRA [blackbox] and compute FSInfo with respect to the token embeddings.
Fig. 3 presents the layer-averaged FSInfo and reconstruction MSE. Overall, FSInfo shows the expected negative correlation with MSE across all datasets: higher FSInfo indicates greater privacy leakage, enabling the DRA adversary to reconstruct inputs more accurately. This aligns with the information-theoretic interpretation of FSInfo—larger FSInfo implies more exposed information and reduced uncertainty for the adversary. However, the correlation is less clear for VGG-9, where the final data point shows an upward trend (Fig. 3(a)). This may indicate that the adversary did not fully exploit the leaked information or that its capability is limited relative to the leakage level. The last two points correspond to the 4th and 2nd layers of VGG-9; the 4th layer produces higher-dimensional smashed data, which likely allows the adversary to capture leaked information more effectively than the 2nd layer.
These evaluations collectively demonstrate that FSInfo remains effective across multiple data types (image, tabular, text), model architectures (FCNs, CNNs, transformers), and application scenarios. Overall, FSInfo is highly consistent with actual attack performance and reliably tracks changes in privacy leakage under different system configurations, supporting its effectiveness as a privacy risk indicator.
V-C Evaluating FSInfo among Defenses
(a) inv_dFIL_def
(b) Nopeek
Figure 4: FSInfo under different defenses and strengths.
As a privacy metric, FSInfo can be used to analyze and monitor privacy risks in DNN-based systems. In this section, we demonstrate its practical relevance by showing that practitioners can also use FSInfo to assess the split inference system employed by defense mechanisms. We examine how privacy leakage changes under different defense strengths using ResNet-18 on CIFAR-10, CelebA, and TinyImageNet (T-ImageNet), with the split point set at the first convolutional layer. We apply FSInfo to SI systems equipped with one regularization-based defense (Nopeek [ICDM-20-nopeek]) and one closed-form defense (inv_dFIL_def [maeng2023bounding]). Defense strength is controlled via hyperparameters: for Nopeek, a larger regularization weight $\alpha$ implies stronger protection, while for inv_dFIL_def, a higher expected MSE lower bound $\frac{1}{dFIL}$ corresponds to stronger defense.
Fig. 4 reports FSInfo values under different defense strengths. The horizontal axis denotes the defense hyperparameters, and the vertical axis corresponds to the datasets. Overall, stronger defenses lead to lower privacy leakage, demonstrating the consistency of FSInfo across datasets and defense mechanisms. An exception occurs on CelebA when the Nopeek hyperparameter increases from 0.5 to 0.7, where privacy leakage slightly increases. This suggests that at low defense strengths, the client model’s optimization may still be dominated by the cross-entropy loss rather than the regularization term.
V-D Evaluating FSInfo by Visual Perception
@@
(a) VGG-5, VGG-9, ResNet-18 on CIFAR-10.
(b) VGG-5 and VGG-9 on MNIST.
Figure 6: SSIM vs. FSInfo on CIFAR-10 and MNIST.
We next examine whether FSInfo aligns with human perception of reconstructed images. Using CIFAR-10, we visualize reconstructions from different layers of ResNet-18 and compute the layer-averaged FSInfo. As shown in Fig. 5, image clarity decreases as FSInfo decreases, indicating a strong correspondence between FSInfo and perceived visual quality. To further validate this, we plot SSIM against FSInfo in Fig. 6. Overall, SSIM positively correlates with FSInfo, suggesting that greater privacy leakage (higher FSInfo) yields reconstructions that are more discernible to humans. The slight downward trend for VGG-9 in Fig. 6(a) likely stems from the same cause noted earlier (Fig. 3(a))—the adversary does not fully exploit the leaked information. Moreover, in Fig. 6(b), the SSIM trend for VGG-9 diverges from that of MSE (Fig. 3(b)), illustrating that empirical attack metrics can sometimes disagree with each other.
V-E Evaluating FSInfoGuard
Figure 7: Task accuracy v.s. MSE on CelebA and ResNet-18.
This section examines whether FSInfo can be used to directly optimize the privacy–utility trade-off (PUT) in practical defense design by comparing FSInfoGuard with existing defenses. In split inference, defensive methods obfuscate the smashed data before transmission to reduce privacy leakage. However, when these defenses rely on privacy metrics that inaccurately estimate leakage, they often introduce excessive noise, leading to significant performance degradation and a poor PUT.
We conduct experiments on CelebA using ResNet-18 and compare FSInfoGuard with Shredder, Nopeek, and inv_dFIL_def. Shredder and inv_dFIL_def derive their Gaussian noise from their respective privacy metrics, whereas Nopeek only optimizes the bottom model without adding noise. Defense strength is controlled by varying each method’s hyperparameters. As shown in Fig. 7, where the $x$-axis indicates reconstruction MSE and the $y$-axis indicates task accuracy, FSInfoGuard achieves the strongest privacy–utility trade-off. This demonstrates that FSInfo provides more effective guidance for designing privacy-preserving mechanisms and accurately quantifies privacy leakage, enabling noise calibration that protects privacy without substantially degrading performance.
V-F The Impact of Intrinsic Data Characteristics